Required information The following Information applies to the questions displayed below.] A manufactured product has the following Information for June Direct eaterials Direct labor (6 lbs. e $7 per lb.) (2 hrs. e $15 per hr.) (2 hrs. e $13 per hr.) 46, 800 lbs. $7.10 per 1b 15, 100 hrs. $15.50 per hr $206, 800 Units sanufactured 7,700 Compute the direct materials price variance and the direct materials quantity varlance. Indicate whether each varlance is favorable or AQ Actual Quantity SQ Standard Quantity AP Actual Price SP- Standard Price Cost AQ AP sQ AP AQ SP Direct labor rate variance Direct materials quantity variance Direct materials price variance
Answer #1

To calculate direct material price and quantity variance:

Actual quantity (AQ) = 46800lbs

Standard quantity (SQ) = 46200lbs (7700units*6lbs)

Actual price (AP) = $7.10 per lb

Standard price (SP) = $7 per lb

Therefore direct material price variance = AQ*(SP-AP)

= 46800*(7-7.10)

=46800* -0.10

= -4680 i.e unfavorable

Direct material quantity variance= SP*(SQ-AQ)

=7*(46200-46800)

=7* -600

= -4200 i.e unfavorable

Total direct material variance = (SP*SQ) - (AP*AQ)

=(7*46200) - (7.10*46800)

= 323400 - 332280

= -8880 i.e unfavorable

Direct labour rate variance =

Actual hours (AH) = 15100 hrs

Standard rate (SR) = $15 per hr

Actual rate (AR) = $15.50 per hr

Direct labour rate variance = AH*(SR-AR)

=15100*(15-15.50)

=15100* -0.50

= -7550 i.e unfavorable